OpenStudy (anonymous):

You owe $1,534.09 on a credit card that has an interest rate of 15.7% APR. You decide to put $250.00 into a savings account earning a 3.5% APR and another $500.00 toward the credit card at the beginning of the month. What is the difference in interest saved at the end of a full month between this decision and if you had paid the entire $750.00 toward the credit card?

OpenStudy (dumbcow):

if you paid entire 750 toward credit card: \[Interest = 750*(.157/12)\] if you put 250 in savings \[Interest = 500*(.157/12)+250(.035/12)\] difference \[= 250*\frac{.157-.035}{12}\]

OpenStudy (anonymous):

-4.02 is that the answer?

OpenStudy (dumbcow):

no

OpenStudy (anonymous):

wold it be 9.81- - 4.02?

OpenStudy (dumbcow):

the 4.02 part is wrong

OpenStudy (anonymous):

what is the interest for that one?

OpenStudy (dumbcow):

7.27

OpenStudy (anonymous):

so it would be 9.81- 7.27

OpenStudy (anonymous):

can you give me the answwer for this question cuz its getting late and imma have to go to bed in a lil bit.

OpenStudy (dumbcow):

"so it would be 9.81- 7.27" this is correct

OpenStudy (anonymous):

but whats the answr to thequestion